OpenStudy (anonymous):

verify that the function satisfies the hypotheses of the mean value theorem on the given interval. Then find all numbers c that satisfy the conclusion of the Mean Value Theorem. f(x)=2x^2-3x+1 [0,2]

OpenStudy (anonymous):

I got 4x-3=3/2 ...

OpenStudy (anonymous):

wait 4x-3=-1

OpenStudy (anonymous):

x=2/4=1/2

OpenStudy (anonymous):

c=1/2